The siren of a fire engine that is driving northward at 30.0m/s emits a sound of frequency 2000Hz . A truck in front of this fire engine is moving northward at 21.0m/s .
Answer #1

f=(c+v_r)f_o/(c+v_s)

So

f=(340.3-21)*2000/(340.3-30)=2058Hz